# Phase engineering of MoS2 monolayers: A pathway to enhanced lithium-polysulfide battery performance

### v1

@@ I Introduction
Li–S batteries operate through a multistep redox mechanism in which elemental sulfur (S8) is progressively reduced to form lithium polysulfides (Li2Sn), including long-chain species (Li2S6, Li2S8) and short-chain intermediates (Li2S2, Li2S4), before ultimately yielding solid Li2S during discharge.Su and Manthiram (2012b) This sequence governs critical aspects of cell performance, such as Coulombic efficiency, self-discharge, and cycle life. Lithium polysulfides (LimSn) thus play a central role in determining the electrochemical behavior of Li–S cells. Figure 1(top) shows the DFT-optimized geometries of key intermediates. Modeling this full polysulfide manifold is essential to reproduce the shuttle effect and to evaluate surface-engineering strategies that suppress it. Li et al. (2018b); Song et al. (2020a)
@@ III.2 Adsorption Properties of LimSn Molecules
We evaluate the adsorption of LimSn molecules on both 2H and 1T MoS2 monolayers using three key descriptors: adsorption energy ($E_{\mathrm{ads}}$), charge transfer ($\Delta Q$), and vertical distance ($h$) between the adsorbate and the surface, as shown in Figure 2. The adsorption energy is defined as:
@@
Figure 2(a) shows that 1T-MoS2 exhibits stronger adsorption energies than 2H-MoS2, especially for smaller polysulfides such as Li2S and Li2S2. In the context of Li-S batteries, this suggests that 1T-MoS2 more effectively anchors polysulfides, suppressing the shuttle effect and improving cycling stability and energy efficiency.Manthiram (2020); Su and Manthiram (2012a)
@@
Overall, the 1T-MoS2 phase exhibits superior performance compared to 2H-MoS2 in terms of adsorption strength, charge donation, and interfacial proximity, indicating a greater ability to anchor polysulfides and facilitate redox reactions. This behavior is consistent with prior studies reporting that metallic TMDs generally interact more strongly with polysulfide species than their semiconducting counterparts.Hojaji et al. (2022); Liu et al. (2021) However, this enhanced interaction can also pose potential limitations: overly strong binding may hinder the desorption of reaction intermediates, particularly during the late stages of the redox cycle, and reduce reversibility under dynamic conditions.Huang et al. (2018); Hojaji et al. (2022) These findings suggest that while 1T domains offer significant advantages for polysulfide immobilization, careful control of their coverage and local environment is necessary to avoid kinetic bottlenecks associated with overbinding. Nevertheless, the weaker interaction of 2H-MoS2 may offer advantages by enabling controlled polysulfide mobility, potentially facilitating redox reversibility while limiting over-anchoring. This tradeoff highlights the value of hybrid systems that combine both phases to optimize sulfur cathode performance in Li-S batteries.
Our DFT results are in agreement with previous experimental reports. For instance, oxygen-doped MoS2 with partial 1T character achieves a discharge capacity of 1015 mAh g-1 with 78.5% retention over 300 cycles,Ren et al. (2024) consistent with enhanced charge transfer in metallic domains. Similarly, 1T-MoS2/MnO2 composites promote a transition from liquid-solid to solid-solid Li2S formation, enabling more uniform nucleation,Qing et al. (2024) while MCNT/MoS2/S hybrids maintain capacities near 460 mAh g-1 after 300 cycles,Guo et al. (2020) although the influence of conductive additives complicates direct interpretation.
@@ III.3 Electronic Properties
By contrast, 1T-MoS2 exhibits pronounced changes in its DOS upon adsorption. Even for S8 [Figure 4(a)], significant hybridization occurs, with enhanced density of states appearing below the Fermi level (EF). This effect intensifies with lithium-containing species: Li2S4 and Li2S6 induce substantial DOS redistribution near EF [Figures 4(d) and (e)], consistent with the stronger charge transfer and binding energies reported in Section III.2. These changes reflect the metallic character of 1T-MoS2 and its ability to accept electrons from LimSn, facilitating improved electronic coupling and redox activity. These distinct responses arise from the inherent electronic differences between the two polymorphs. The semiconducting 2H-MoS2 shows minimal perturbation upon adsorption, with only weakly localized states near the band edges. In contrast, the metallic 1T phase undergoes substantial electronic restructuring, indicating stronger adsorbate-substrate interactions. These findings agree with previous studies reporting enhanced electron accommodation and catalytic activity in metallic TMDs.Manthiram (2020)
From a practical perspective, combining both polymorphs offers a promising strategy: 1T-MoS2 provides high conductivity and abundant active sites for redox reactions, while 2H-MoS2 may help stabilize the structure and suppress undesired electron leakage. This concept of phase complementarity is increasingly adopted in Li-S cathode design, leveraging the synergy between conductive and semiconducting domains to achieve both efficient charge transport and effective sulfur confinement.Lukowski et al. (2013)
@@ III.4 Delithiation Pathway
Figure 6 complements the NEB results by depicting the fully relaxed atomic configurations and corresponding adsorption energies along the delithiation sequence on 1T-MoS2. Throughout the process-from Li2S4 to LiS4 to S4-the system maintains strong interfacial coupling, accompanied by a progressive increase in charge transfer ($\Delta Q$). Specifically, $\Delta Q$ increases from 0.25 e- in the initial state to 0.36 e- for LiS4, and further to 0.44 e- for the final S4 product. This monotonic growth in $\Delta Q$ reflects efficient charge accommodation by the metallic substrate, enabling stabilization of partially oxidized intermediates via delocalized electronic states. The smooth energy landscape and persistent binding interactions suggest a continuous and energetically favorable reaction path that supports rapid redox kinetics and effective sulfur utilization. These properties, combined with reduced reaction barriers, reinforce the role of 1T-MoS2 as a highly effective anchoring and catalytic platform for lithium polysulfide conversion in Li-S batteries.Manthiram (2020)
@@ III.5 Comparison 1T-MoS2 Polymorph
Therefore, 1T-MoS2 offers a more balanced profile: it exhibits sufficient binding to suppress polysulfide migration while maintaining low kinetic barriers to sustain redox activity. This intermediate behavior avoids the desorption bottlenecks of overbinding observed in vacancy-rich 1T phases, suggesting that 1T may serve as a more versatile and tunable platform for Li–S battery cathodes.
@@ III.6 Thermochemistry (Gas-phase approach)
The thermodynamic stability of adsorbed polysulfides depends not only on their binding strength at 0 K, but also on the temperature and the chemical environment-particularly the presence of solvent. These factors are critical for real operating conditions in Li-S batteries, where temperature variations and electrolyte composition influence desorption, shuttling, and redox reversibility. To capture these effects, we compute Gibbs free energy changes ($\Delta G$) for LimSn adsorption on 1T-MoS2 across a range of temperatures and under both gas-phase and solvated conditions. Within this framework, we compute the Gibbs free energy change for adsorption as:
@@
From a design perspective, electrolyte formulation must balance two competing effects: interactions strong enough to immobilize polysulfides and suppress the shuttle effect yet sufficiently weak to avoid overbinding that could hinder redox kinetics. Although DMC solvation reduces adsorption strength, it may enhance molecular mobility and help prevent electrode passivation. Solvation also affects the temperature threshold at which adsorption becomes thermodynamically unfavorable. Our calculations place this crossover near 473 K in the gas phase and approximately 393 K under implicit solvation, showing the role of the solvent in tuning the thermal stability and operational regimes of adsorbed polysulfide species.Manthiram (2020); Su and Manthiram (2012a)
@@
III.7 Spectral Analysis:
Infrared (IR) spectroscopy provides a direct and experimentally accessible probe of molecular bonding environments and interfacial interactions. To complement the thermodynamic analysis, Figure 7 presents the calculated IR spectra of Li2S4 adsorbed on a 1T-MoS2 cluster, evaluated under both gas-phase and solvated conditions. These spectra, obtained via DFT calculations using ORCA, reveal substantial vibrational shifts upon adsorption, indicative of strong chemical coupling between the polysulfide and the metallic surface.
@@ III.7 Spectral Analysis:
Significantly, adsorption on 1T-MoS2 consistently induces downshifts in several Li–S modes, often associated with bond elongation or weakening due to interfacial charge transfer and substrate-induced strain.Zhao and Truhlar (2008) These vibrational signatures correlate with the increase in charge transfer ($\Delta Q$) and stronger adsorption energies ($E_{\mathrm{ads}}$), reinforcing the role of 1T-MoS2 as an electronically responsive and structurally adaptive host for polysulfide immobilization.
@@ IV Final Remarks
This study employs first-principles calculations to evaluate the potential of two MoS2 polymorphs, 2H and 1T, as anchoring materials for lithium-sulfur (Li-S) battery cathodes. We assess their ability to suppress polysulfide shuttling and enhance redox kinetics through detailed analysis of their electronic structure, adsorption energetics, and thermochemical response. Our results demonstrate that the metallic 1T phase consistently outperforms the semiconducting 2H phase in adsorption strength, charge transfer, and electronic coupling. These properties enable stronger immobilization of lithium polysulfides and facilitate faster redox dynamics. Conversely, the 2H phase, despite its weaker interactions, may improve mechanical integrity and reduce parasitic conductivity, suggesting a complementary role in hybrid architectures. Thermodynamic calculations reveal that solvation substantially weakens adsorption, lowering the polysulfide desorption threshold by approximately 80 K. This trend underscores the importance of considering solvent effects when designing interfacial chemistries. Simulated infrared spectra further confirm strong chemical coupling between Li2S4 and the 1T surface, manifested as pronounced redshifts in vibrational modes, consistent with significant charge transfer and interfacial restructuring.
@@
From a materials design perspective, our results suggest that optimizing the coverage of 1T domains can enhance polysulfide anchoring, but excessive metallic character may lead to overbinding and hinder desorptionHuang et al. (2018); Hojaji et al. (2022). Operational stability may be improved by maintaining temperatures below the calculated desorption crossover (393 K under solvation), while targeted phase or defect engineering, through doping, strain, or electrochemical control, can be employed to tune the 2H/1T ratio and activate interphase regions. Additionally, electrolyte formulation should reduce polysulfide solubility while preserving sufficient mobility to avoid passivation or irreversible binding.
@@
This work highlights the promise of phase-engineered MoS2 as a modular platform for next-generation lithium-polysulfide cathodes. A synergistic approach integrating computational screening with targeted experimental efforts will be key to realizing durable, high-performance, and tunable energy storage systems.

### v2

@@ I Introduction
Li–S batteries operate through a multistep redox mechanism in which elemental sulfur (S8) is progressively reduced to form lithium polysulfides (Li2Sn), including long-chain species (Li2S6, Li2S8) and short-chain intermediates (Li2S2, Li2S4), before ultimately yielding solid Li2S during discharge.Su and Manthiram (2012b) This sequence governs critical aspects of cell performance, such as Coulombic efficiency, self-discharge, and cycle life. Lithium polysulfides (LimSn) thus play a central role in determining the electrochemical behavior of Li–S cells. Figure 1(top) shows the DFT-optimized geometries of key intermediates. Modeling the full polysulfide manifold is essential to reproduce the shuttle effect and to evaluate surface-engineering strategies that suppress it;Li et al. (2018b); Song et al. (2020a) in this work, we address this challenge by focusing on a representative species (Li2S4), which captures key features of the intermediate redox steps.
@@ III.2 Adsorption Properties of LimSn Molecules
We evaluate the adsorption of LimSn molecules on both 2H and 1T MoS2 monolayers by examining three key descriptors: adsorption energy ($E_{\mathrm{ads}}$), charge transfer ($\Delta Q$), and vertical distance ($h$) between the adsorbate and the surface, as shown in Figure 2. The adsorption energy is defined as follows:
@@
Figure 2(a) shows that 1T-MoS2 exhibits stronger adsorption energies than 2H-MoS2, especially for smaller polysulfides such as Li2S and Li2S2. In the context of Li–S batteries, this suggests that 1T-MoS2 more effectively anchors polysulfides, suppressing the shuttle effect and improving cycling stability and energy efficiency.Manthiram (2020); Su and Manthiram (2012a)
@@
Overall, the 1T-MoS2 phase exhibits superior performance compared to 2H-MoS2 in terms of adsorption strength, charge donation, and interfacial proximity, indicating a greater ability to anchor polysulfides and facilitate redox reactions. This behavior is consistent with prior studies reporting that metallic TMDs generally interact more strongly with polysulfide species than their semiconducting counterparts.Hojaji et al. (2022); Liu et al. (2021) However, this enhanced interaction can also pose potential limitations: overly strong binding may hinder the desorption of reaction intermediates, particularly during the late stages of the redox cycle, and reduce reversibility under dynamic conditions.Huang et al. (2018); Hojaji et al. (2022) These findings suggest that while 1T domains offer significant advantages for polysulfide immobilization, careful control of their coverage and local environment is necessary to avoid kinetic bottlenecks associated with overbinding. Nevertheless, the weaker interaction of 2H-MoS2 may offer advantages by enabling controlled polysulfide mobility, potentially facilitating redox reversibility while limiting over-anchoring. This tradeoff highlights the value of hybrid systems that combine both phases to optimize sulfur cathode performance in Li–S batteries.
Our DFT results are in agreement with previous experimental reports. For instance, oxygen-doped MoS2 with partial 1T character achieves a discharge capacity of 1015 mAh g-1 with 78.5% retention over 300 cycles,Ren et al. (2024) consistent with enhanced charge transfer in metallic domains. Similarly, 1T-MoS2/MnO2 composites promote a transition from liquid-solid to solid-solid Li2S formation, enabling more uniform nucleation.Qing et al. (2024) Additionally, MCNT/MoS2/S hybrids maintain capacities near 460 mAh g-1 after 300 cycles,Guo et al. (2020) although the influence of conductive additives complicates direct interpretation.
@@ III.3 Electronic Properties
By contrast, 1T-MoS2 phase exhibits significant changes in its DOS upon adsorption. Even for S8 [Figure 4(a)], significant hybridization occurs, with enhanced density of states appearing below the Fermi level (EF). This effect intensifies with lithium-containing species: Li2S4 and Li2S6 induce substantial DOS redistribution near EF [Figures 4(d) and (e)], consistent with the stronger charge transfer and binding energies reported in Section III.2. These changes reflect the metallic character of 1T-MoS2 and its ability to accept electrons from LimSn, facilitating improved electronic coupling and redox activity. These distinct responses arise from the inherent electronic differences between the two polymorphs. The semiconducting 2H-MoS2 shows minimal perturbation upon adsorption, with only weakly localized states near the band edges. In contrast, the metallic 1T phase undergoes substantial electronic restructuring, indicating stronger adsorbate-substrate interactions. These findings agree with previous studies that report improved electron accommodation and catalytic activity in metallic TMDs.Manthiram (2020)
From a practical perspective, combining both polymorphs offers a promising strategy: 1T-MoS2 provides high conductivity and abundant active sites for redox reactions, while 2H-MoS2 may help stabilize the structure and suppress undesired electron leakage. This concept of phase complementarity is increasingly adopted in Li–S cathode design, leveraging the synergy between conductive and semiconducting domains to achieve both efficient charge transport and effective sulfur confinement.Lukowski et al. (2013)
@@ III.4 Delithiation Pathway
Figure 6 complements the NEB results by depicting the fully relaxed atomic configurations and corresponding adsorption energies along the delithiation sequence on 1T-MoS2. Throughout the process-from Li2S4 to LiS4 to S4-the system maintains strong interfacial coupling, accompanied by a progressive increase in charge transfer ($\Delta Q$). Specifically, $\Delta Q$ increases from 0.25 e- in the initial state to 0.36 e- for LiS4, and further to 0.44 e- for the final S4 product. This monotonic growth in $\Delta Q$ reflects efficient charge accommodation by the metallic substrate, enabling stabilization of partially oxidized intermediates via delocalized electronic states. The smooth energy landscape and persistent binding interactions suggest a continuous and energetically favorable reaction path that supports rapid redox kinetics and effective sulfur utilization. These properties, combined with reduced reaction barriers, reinforce the role of 1T-MoS2 as a highly effective anchoring and catalytic platform for lithium polysulfide conversion in Li–S batteries.Manthiram (2020)
@@ III.5 Comparison 1T-MoS2 Polymorph
Therefore, 1T-MoS2 structure offers a more balanced profile: it exhibits sufficient binding to suppress polysulfide migration while maintaining low kinetic barriers to sustain redox activity. This intermediate behavior avoids the desorption bottlenecks of overbinding observed in vacancy-rich 1T phases, suggesting that 1T may serve as a more versatile and tunable platform for Li–S battery cathodes.
@@ III.6 Thermochemistry (Gas-phase approach)
The thermodynamic stability of adsorbed polysulfides depends not only on their binding strength at 0 K, but also on the temperature and the chemical environment-particularly the presence of solvent. These factors are critical for real operating conditions in Li–S batteries, where temperature variations and electrolyte composition influence desorption, shuttling, and redox reversibility. To capture these effects, we compute Gibbs free energy changes ($\Delta G$) for LimSn adsorption on 1T-MoS2 across a range of temperatures and under both gas-phase and solvated conditions. Within this framework, we compute the Gibbs free energy change for adsorption as:
@@
From a design perspective, electrolyte formulating electrolytes requires balancing two competing effects: interactions strong enough to immobilize polysulfides and suppress the shuttle effect yet sufficiently weak to avoid overbinding that could hinder redox kinetics. Although DMC solvation reduces adsorption strength, it may enhance molecular mobility and help prevent electrode passivation. Solvation also affects the temperature threshold at which adsorption becomes thermodynamically unfavorable. Our calculations place this crossover near 473 K in the gas phase and approximately 393 K under implicit solvation, showing the role of the solvent in tuning the thermal stability and operational regimes of adsorbed polysulfide species.Manthiram (2020); Su and Manthiram (2012a)
@@
III.7 Spectral Analysis
Infrared (IR) spectroscopy serves as a direct and experimentally accessible probe of molecular bonding environments and interfacial interactions. To complement the thermodynamic analysis, Figure 7 displays the IR spectra of Li2S4 adsorbed on a 1T-MoS2 cluster, computed under both gas-phase and solvated conditions. The spectra, obtained from DFT calculations using ORCA, exhibit pronounced vibrational shifts upon adsorption, reflecting strong chemical coupling between the polysulfide species and the metallic surface.
@@ III.7 Spectral Analysis
Significantly, adsorption on 1T-MoS2 consistently induces downshifts in several Li–S modes, often associated with bond elongation or weakening due to interfacial charge transfer and substrate-induced strain.Zhao and Truhlar (2008) These vibrational signatures correlate with the increase in charge transfer ($\Delta Q$) and stronger adsorption energies ($E_{\mathrm{ads}}$), reinforcing the role of 1T-MoS2 as an electronically responsive and structurally adaptive host for polysulfide immobilization. Importantly, the predicted spectral features offer clear experimental targets for verification. Infrared and Raman spectroscopy can directly probe the Li–S stretching and bending modes and monitor their redshifts upon adsorption on MoS2 surfaces. By tracking the emergence of new vibrational features associated with Mo-S-Li interfacial interactions and following the evolution of these peaks during battery cycling, experimental studies can sensitively assess the anchoring behavior, catalytic activity, and phase identity predicted by our calculations.
@@ IV Final Remarks
This study employs first-principles calculations to evaluate the potential of two MoS2 polymorphs, 2H and 1T, as anchoring materials for lithium-sulfur (Li–S) battery cathodes. We assess their ability to suppress polysulfide shuttling and enhance redox kinetics through detailed analysis of their electronic structure, adsorption energetics, and thermochemical responses. Our results demonstrate that the metallic 1T phase consistently outperforms the semiconducting 2H phase in adsorption strength, charge transfer, and electronic coupling. These properties enable stronger immobilization of lithium polysulfides and facilitate faster redox dynamics. Conversely, the 2H phase, despite its weaker interactions, may improve mechanical integrity and reduce parasitic conductivity, suggesting a complementary role in hybrid architectures. Thermodynamic calculations reveal that solvation substantially weakens adsorption, lowering the polysulfide desorption threshold by approximately 80 K. This trend underscores the importance of considering solvent effects when designing interfacial chemistries. Simulated infrared spectra further confirm strong chemical coupling between Li2S4 and the 1T surface, manifested as pronounced redshifts in vibrational modes, consistent with significant charge transfer and interfacial restructuring. These predicted vibrational signatures provide clear experimental targets: infrared and Raman spectroscopy can directly probe the Li–S stretching and bending modes, track their redshifts upon adsorption, and monitor the emergence of new Mo–S–Li interfacial features during battery cycling, offering a sensitive route to validate the proposed anchoring and catalytic behavior.
@@
From a materials design perspective, our results suggest that optimizing the coverage of 1T domains can enhance polysulfide anchoring, but excessive metallic character may lead to overbinding and hinder desorptionHuang et al. (2018); Hojaji et al. (2022). Operational stability may be improved by maintaining temperatures below the calculated desorption crossover (393 K under solvation), while targeted phase or defect engineering, through doping, strain, or electrochemical control, can be employed to tune the 2H/1T ratio and activate interphase regions. Specifically, chemical exfoliation, alkali metal intercalation, electron beam irradiation, and lithium intercalation under controlled potentials offer promising strategies to trigger the 2H to 1T transition, providing tunable pathways to adjust the phase composition and optimize the balance between conductivity, stability, and catalytic performance. Additionally, electrolyte formulation should reduce polysulfide solubility while preserving sufficient mobility to avoid passivation or irreversible binding.
Beyond the specific case of Li–S batteries, the phase engineering principles identified here are broadly applicable to other alkali-metal-chalcogen battery systems, such as Na–S, K–S, and Li–Se chemistries. These systems share similar challenges related to shuttle effects, low electronic conductivity, and sluggish redox kinetics. Controlling the relative abundance of 2H and 1T phases in MoS2 or analogous two-dimensional materials offers a general strategy to tune interfacial interactions, enhance catalytic activity, and improve electrochemical performance across a wide range of emerging energy storage technologies. This phase-tuning approach could be further extended to hybrid solid-state architectures and multifunctional electrode designs, opening avenues for cross-platform improvements beyond the Li–S paradigm.
@@
This work highlights the promise of phase-engineered MoS2 as a modular platform for next-generation lithium-polysulfide cathodes. A synergistic approach integrating computational screening with targeted experimental efforts will be key to realizing durable, high-performance, and tunable energy storage systems. Taken together, these insights lay a solid foundation for the rational design of scalable, high-efficiency battery architectures and position phase-engineered two-dimensional materials as essential building blocks for advancing the frontiers of energy storage technology.